OpenStudy (anonymous):

HELP! MEDAL :D The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 15x - 12 g(x) = 3x -2 0.11 -1 -27 0.33 0 -12 1 2 Complete the chart and determine the solution of the equation f(x) = g(x) x = -1 x = 0 x = 1 x = 3

OpenStudy (anonymous):

@Hero @UnkleRhaukus @dan815 @cwrw238

OpenStudy (anonymous):

Better graph http://postimg.org/image/n657ibpb5/

OpenStudy (anonymous):

It just wants you to plug in each value for x in each equation. And then determine what x equals when the two graphs equal the same value. :) I'll do the first one, x=-2 for f(x) as an example. f(x)=15(-2)-12 f(x)=-30-12 f(x)-42 And an example for g(x), when x=0 g(x)=3^0 g(x)=1 Lemme know when you get all the values if you want me to help with the f(x)=g(x) part :)

OpenStudy (anonymous):

Yep I kinda need help..

OpenStudy (anonymous):

Oh with the second part? Lol have you got all the values in the table then?

OpenStudy (anonymous):

Still need kinda help

OpenStudy (fatimab2000):

x=1